OpenStudy (anonymous):

How do you find the relative extrema of (x^3)-(3x^2)+3x-1

OpenStudy (anonymous):

take the derivative.....equate it to zero 3x^2-6x+3=0 x^2-2x+1=0 (x-1)^2=0 x=1

OpenStudy (anonymous):

Really? I got the exact same thing you did and I had figured something was off just because it seemed too easy >_> Thanks a bunch though.